OpenStudy (barreraa):

What is the relationship between the lines determined by these two equations? 3x + 6y = 8 y = 2x – 8 (Points : 4) parallel perpendicular neither parallel nor perpendicular they are the same line

OpenStudy (anonymous):

\[3x + 6y = 8\] \[6y =-3x+ 8\] \[y =-3/6x+ 8/6\] \[y =-1/2x+ 4/3\]

OpenStudy (anonymous):

slope =-1/2

OpenStudy (anonymous):

y=2x-8 slope =2

OpenStudy (anonymous):

slope 1*slope 2=\[2*\frac{ -1 }{ 2 }=-1\]

OpenStudy (anonymous):

perpendicular